OpenStudy (anonymous):

factor completely m^2+7m+10 y^2-13+06 x^2-8x+7 x^2=9x+18 r^2+4r+3 n^2-3n+2 x^2-2x+1 k^2-16k+18 x^2+3x-4 m^2-2m-8 d^2+6d-40 p^2-15p-54 k^2-8k-9 y^2-18y+45

OpenStudy (abb0t):

Whoa! That's a lot of questions. Lol.

OpenStudy (anonymous):

to factor those you must find their roots, that is for example for what x => x^2+3x-4=0. So you must solve a lot of quadratic equations...

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

so you don't know how to solve quadratic equations? or you just someone else to do this for you?

OpenStudy (abb0t):

I'm not going to do your homework for you. That won't help you if I do it for you, but i will guide you so you can work on it on your own. \[ax^2+bx+c\] What you want to find is basically two numbers when multiplied give you C, but, when you add those two numbers, you get "b". Does that make sense? I will do the first one for you: \[m^2+7m+10\] notice that it's in the form: \[ax^2+bx+c\] where b = 7 and c = 10 find those two numbers when multiplied (5 x 2) = 10 but when added (5 + 2) = 7 As you ca see, those two numbers are 5 and 2. So you factor it out: \[m^2+5m+2m+10 = m(m+5) + 2(m+5) = (m+2)(m+5)\] You can check to make sure that those two numbers give you the original polynomial by doing the whole FOIL method thing if you are unsure.

OpenStudy (anonymous):

k thank you

OpenStudy (abb0t):

However, if there does come a problem where you cannot find two numbers when multiplied = c and added give you "b" then you must use the quadratic formula: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

and you are sure about this ?

OpenStudy (abb0t):

Do your homework!

OpenStudy (abb0t):

daddy*

OpenStudy (anonymous):

Aha Wow